# EPW varieties as moduli spaces on ordinary GM surfaces and special GM threefolds

### v2

@@ 3.1. Stability conditions on triangulated categories
To reconstruct EPW varieties, we need to use the notion of stability conditions introduced by Bridgeland [10]. Here we will adapt a more recent formulation of stability conditions proposed in [6, Appendix 1] based on the locally-finite full stability conditions in [10, 11].
@@
  • one has $Z(\textup{{v}}(E))\in\mathbb{R}_{>0}\cdot\exp(\pi\phi\sqrt{-1})$ for any $0\neq A\in\mathcal{P}(\phi)$;
@@
  • there exists a quadratic form $Q$ on the vector space $\Lambda_{\mathbb{R}}$ such that $\ker(Z_{\mathbb{R}})$ is negative definite with respect to $Q$, and one has $Q(\textup{{v}}(E))\geq 0$ for any $A\in\mathcal{P}(\phi)$ and any $\phi\in\mathbb{R}$.
@@
Let $\sigma$ be stability condition, then the Harder–Narasimhan filtration for an object $E$ is unique up to isomorphism. In this case, the objects $A_{i}$ are called the factors of $E$. Two objects are called $S$-equivalence once they admitting the same Harder–Narasimhan factors.
@@
Let $\sigma=(\mathcal{P},Z)$ be stability condition, then any abelian category $\mathcal{P}(\phi)$ has finite length (see for example [10, page 330]) and an object in $\mathcal{P}(\phi)$ therefore admits a Jordan–Hölder filtration with uniquely determined Jordan–Hölder factors.
@@
Also, there are two mutually-commutative group actions on the complex manifold $\operatorname{Stab}(\mathcal{D})$ due to [10, Lemma 8.2] as following.
@@ 3.2. Moduli spaces of semistable objects
Such a good moduli space $\textsf{{M}}_{\sigma}(v)$ is an algebraic space and the (closed) points in $\textsf{{M}}_{\sigma}(v)$ are $S$-equivalent classes of semistable objects with phase ranging in (1,1](-1,1].
@@
Proposition 3.8 ([11, Section 9] and [4, Proposition 2.3]).
Given $v\in\textsf{{N}}(\mathcal{D})$, then there exists a locally finite set of walls (real codimension one submanifolds with boundary) in the complex manifold $\operatorname{Stab}(\mathcal{D})$, depending only on $v$, such that:
  • The sets of $\sigma$-semistable and $\sigma$-stable objects of class $v$ remain unchanged for any stability condition $\sigma$ in the same chamber;
@@
In general, the classification of walls are not available but one always has the following two situations for a stability condition.
@@
The walls and wall-crossing behaviors of the moduli space $\textsf{{M}}_{\sigma}(v)$ are only well-understood for some stability conditions on K3 surfaces [4, 5, 35] and Enriques surfaces [36]. Since a strongly smooth Gushel–Mukai surface is in particular a K3 surface, the article uses some results regarding walls and wall-crossing for K3 surfaces which are recollected in what follows.
@@ 3.3. Some stability conditions on K3 surfaces
$\mathcal{A}_{\omega,\beta}=\{E\in{\textbf{{D}}}^{b}(S)\,|\,H^{i}(E)=0\textup{ for }i\notin\{-1,0\},H^{-1}(E)\in\mathcal{F}_{\omega,\beta},H^{0}(E)\in\mathcal{T}_{\omega,\beta}\}$
@@
Proposition 3.12 ([11, Lemma 6.2 and Section 7]).
@@
using the above criterion and Proposition 2.4.
@@ 3.4. The walls and wall-crossing on K3 surfaces
Let $S$ be a K3 surface and $v\in\textsf{{H}}(S,\mathbb{Z})$ be a primitive Mukai vector with $\langle v,v\rangle\geq-2$. Suppose that $W\subset\operatorname{Stab}^{{\dagger}}(S)$ be a wall respect to $v$ in the sense of Proposition 3.8 and $\sigma_{0}$ is a generic point on $W$, then for any two $v$-generic stability conditions $\sigma_{\pm}$ in the adjoint chambers with respect to $W$ one has a birational map
@@
and two contractions $\textsf{{M}}_{\sigma_{\pm}}(v)\rightarrow\bar{M}_{\pm}$ contracting $S$-equivalent objects by [4, Theorem 1.4].
@@
Proposition 3.15 ([5, Proposition 5.1]).
Under the setups in this subsection, the Mukai vector of any HN factor of an object $A$ in $\textsf{{M}}_{\sigma_{+}}(v)$ with respect to $\sigma_{-}$ is contained in $\textbf{{H}}_{W}$, and vice versa. The same holds for any object in $\textsf{{M}}_{\sigma_{0}}(v)$. Similarly, any Jordan–Hölder factor of an object in $A\in\textsf{{M}}^{\textsf{{st}}}_{\sigma_{0}}(w)$ for some $w\in\textbf{{H}}_{W}$ will have Mukai vector contained in $\textbf{{H}}_{W}$.
In this case, using the existence of certain semistable objects, one can determine the type of the wall $W$ according to specific vectors in $\textbf{{H}}_{W}$ by [5, Theorem 5.7] when $\langle v,v\rangle>0$.
@@
Let $\langle v,v\rangle>0$, then the wall $W$ is called a fake wall once there are no curves in $\textsf{{M}}_{\sigma_{\pm}}(v)$ that are $S$-equivalent to each other respect to $\sigma_{0}$; a flopping wall once $\bar{M}_{+}=\bar{M}_{-}$ and the induced birational map $\textsf{{M}}_{\sigma_{-}}(v)\dashrightarrow\textsf{{M}}_{\sigma_{+}}(v)$ induces a flopping contraction; a divisorial wall once the induced contractions $\textsf{{M}}_{\sigma_{\pm}}(v)\rightarrow\bar{M}_{\pm}$ are divisorial contractions.
@@
In this case, according to [5, Proposition 9.4 and Section 14], the Jordan–Hölder factors for a strictly semistable object $A$ in $\textsf{{M}}_{\sigma_{0}}(v)$ are spherical $\sigma_{\pm}$-stable objects. Moreover, one notices that the vector $w^{\prime}=v-w$ also satisfies $\langle w^{\prime},w^{\prime}\rangle=-2$ and $\langle v,w^{\prime}\rangle=1$ in the setting of Theorem 3.18, so a general flopping wall is induced by two spherical stable objects.
@@ 4. The passage to special Gushel–Mukai threefolds
4.1. Group actions and equivariant categories
We fix a finite group $G$ and a category $\mathcal{D}$.
Definition 4.1.
A (right) action of $G$ on $\mathcal{D}$ is an assignment ${\textup{{{A}}}}_{(-)}\colon G\rightarrow\operatorname{Aut}(\mathcal{D})$ and a set of natural transformations $\{\epsilon_{g,h}\colon{\textup{{{A}}}}_{gh}\stackrel{{\scriptstyle\sim}}{{\rightarrow}}{\textup{{{A}}}}_{h}\circ{\textup{{{A}}}}_{g}\,|\,g,h\in G\}$ such that
${{\textup{{{A}}}}_{g_{1}g_{2}g_{3}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\epsilon_{g_{1},g_{2}g_{3}}}$$\scriptstyle{\epsilon_{g_{1}g_{2},g_{3}}}$${{\textup{{{A}}}}_{g_{3}}\circ{\textup{{{A}}}}_{g_{1}g_{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\textup{{{A}}}}_{g_{3}}\epsilon_{g_{1},g_{2}}}$${{\textup{{{A}}}}_{g_{2}g_{3}}\circ{\textup{{{A}}}}_{g_{1}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\epsilon_{g_{2},g_{3}}{\textup{{{A}}}}_{g_{1}}}$${{\textup{{{A}}}}_{g_{3}}\circ{\textup{{{A}}}}_{g_{2}}\circ{\textup{{{A}}}}_{g_{1}}}$
commutes for all $g_{1},g_{2},g_{3}\in G$.
Given an action of $G$ on $\mathcal{D}$, one can always construct the equivariant category $\mathcal{D}_{G}$ with respect to the $G$-action on $\mathcal{D}$ whose objects and morphisms are respectively the $G$-equivariant objects and $G$-equivariant morphisms.
Definition 4.2.
A $G$-equivariant object of $\mathcal{D}$ is a pair $(A,\phi)$, where $A$ is an object of $\mathcal{D}$ and $\phi=\{\phi_{g}\}_{g\in G}$ is a collection of isomorphisms $\phi_{g}:A\xrightarrow{\sim}{\textup{{{A}}}}_{g}A$ for $g\in G$, such that the diagram
[…]
commutes for all $g,h\in G$. The equivariant structure $\phi$ is often omitted from the notation.
Definition 4.3.
A $G$-equivariant morphism $(A,\phi^{A})\rightarrow(B,\phi^{B})$ of $G$-equivariant objects of $\mathcal{D}$ is a morphism $f\colon A\rightarrow B$ in $\mathcal{D}$ such that the diagram
[…]
commutes for every $g\in G$.
Suppose now that $G$ is abelian, then one can define a natural action of the dual group $G^{\vee}:=\operatorname{Hom}(G,\mathbb{C}^{\times})$ on $\mathcal{D}_{G}$ from a $G$-action on $\mathcal{D}$: for $\chi\in G^{\vee}$ and $(A,\phi)\in\mathcal{D}$ one sets
$\psi_{\chi}(A,\{\phi_{g}\}_{g\in G})=(A,\{\phi_{g}\}_{g\in G})\otimes\chi=(A,\{\phi_{g}\cdot\chi(g)\})$
The equivariant objects $\psi_{\chi_{1}}(\psi_{\chi_{2}}(A,\phi))$ and $\psi_{\chi_{2}\chi_{1}}(A,\phi)$ are the same, so one can just let the required natural transformations $\psi_{\chi_{1}}\circ\psi_{\chi_{2}}\rightarrow\psi_{\chi_{2}\chi_{1}}$ be identities.
Theorem 4.4 ([19, Theorem 4.2]).
Consider a finite abelian group $G$ acting on an idempotent complete category $\mathcal{D}$, then there exists an equivalence
${\textup{{{X}}}}\colon(\mathcal{D}_{G})_{G^{\vee}}\cong\mathcal{D}$
such that the $G$-action is identified with the $(G^{\vee})^{\vee}$-action by $(G^{\vee})^{\vee}\cong G$.
@@ 4.1. Group actions and equivariant categories
given on objects by sending $A$ to $\bigoplus_{g\in G}{\textup{{{A}}}}_{g}A$ with the natural $G$-equivariant structure (one can see, for example, [19, Lemma 3.8]). Moreover, one has the following observation:
@@
are exact. Then, due to Proposition 4.5, the equivalence X in Theorem 4.4 respects exact triangles and therefore an equivalence of triangulated categories.
@@ 4.2. Equivariant stability conditions and covering maps
Proposition 4.7.
In the setup of Theorem 4.6, the stability condition $(\sigma_{\mathfrak{S}_{2}})_{\mathfrak{S}_{2}^{\vee}}$ corresponds to the stability condition $2\cdot\sigma=(\mathcal{P},2Z)$ under the equivalence ${\textup{{{X}}}}\colon(\mathcal{D}_{\mathfrak{S}_{2}})_{\mathfrak{S}_{2}^{\vee}}\cong\mathcal{D}$ of Theorem $4.$
@@ 4.3. The Kuznetsov components and the involutions
The numerical Grothendieck group $\textsf{{N}}({\textbf{{Ku}}}(X))$ has rank 2. Denotes the first Chern class of the polarization $\mathscr{O}_{X}(1)$ by $h$, then one has a basis $\kappa_{1},\kappa_{2}$ of $\textsf{{N}}({\textbf{{Ku}}}(X))$ with
$\textup{{ch}}(\kappa_{1})=1-\frac{1}{5}h^{2}\quad\textup{and}\quad\textup{{ch}}(\kappa_{1})=2-h+\frac{1}{12}h^{3}$
@@
The opposite of $X$ is a strongly smooth Gushel–Mukai surface $S$ and one has a covering involution $f\colon X\rightarrow X$ for the double cover $X\rightarrow M_{S}$ branched over $S\subset M_{S}$. The direct image of the covering involution restricts to an involution on the Kuznetsov component ${\textbf{{Ku}}}(X)$ and induces a $\mathfrak{S}_{2}$-action on ${\textbf{{Ku}}}(X)$. According to [30, 43], the equivariant category ${\textbf{{Ku}}}(X)_{\mathfrak{S}_{2}}$ is equivalent to ${\textbf{{D}}}^{b}(S)$ and the dual $\mathfrak{S}_{2}$-group action on the equivariant category is induced by an involution on ${\textbf{{D}}}^{b}(S)$ say $\Pi$. The involution $\Pi$ induces a Hodge isometry on $\textsf{{H}}(S,\mathbb{Z})$ whose invariant lattice is spanned by $v_{1}=(1,0,-1)$ and $v_{2}=(2,-H,2)$.
@@
$\textup{{inf}}(a\kappa_{1}+b\kappa_{2})$ $=-a(1,0,-1)-b(2,-H,2)$
@@
5. Double duel EPW sextics as Bridgeland moduli spaces

### v3

@@ 3.1. Stability conditions on triangulated categories
To reconstruct EPW varieties, we need to use the notion of stability conditions introduced by Bridgeland [13]. Here we will adapt a more recent formulation of stability conditions proposed in [8, Appendix 1] based on the locally-finite full stability conditions in [13, 14].
@@
  • one has $Z(\textup{{v}}(E))\in\mathbb{R}_{>0}\cdot\exp(\pi\phi\sqrt{-1})$ for any $0\neq E\in\mathcal{P}(\phi)$;
@@
  • there exists a quadratic form $Q$ on the vector space $\Lambda_{\mathbb{R}}$ such that $\ker(Z_{\mathbb{R}})$ is negative definite with respect to $Q$, and one has $Q(\textup{{v}}(E))\geq 0$ for any $E\in\mathcal{P}(\phi)$ and any $\phi\in\mathbb{R}$.
@@
Let $\sigma$ be stability condition, then the Harder–Narasimhan filtration for an object $E$ is unique up to isomorphism. In this case, the objects $A_{i}$ are called the factors of $E$. Two objects are called $S$-equivalence once they have the same Harder–Narasimhan factors.
@@
Let $\sigma=(\mathcal{P},Z)$ be stability condition, then any abelian category $\mathcal{P}(\phi)$ has finite length (see for example [13, page 330]) and an object in $\mathcal{P}(\phi)$ therefore admits a Jordan–Hölder filtration with uniquely determined Jordan–Hölder factors.
@@
Also, there are two mutually-commutative group actions on the complex manifold $\operatorname{Stab}(\mathcal{D})$ due to [13, Lemma 8.2] as following.
@@ 3.2. Moduli spaces of semistable objects
Such a good moduli space $\textsf{{M}}_{\sigma}(v)$ is an algebraic space and the (closed) points in $\textsf{{M}}_{\sigma}(v)$ are $S$-equivalence classes of semistable objects with phase ranging in (1,1](-1,1].
@@
Proposition 3.8 ([14, Section 9] and [5, Proposition 3.3]).
Given a primitive $v\in\textsf{{N}}(\mathcal{D})$, then there exists a locally finite set of walls (real codimension one submanifolds with boundary) in the complex manifold $\operatorname{Stab}(\mathcal{D})$, depending only on $v$, such that:
  • The sets of $\sigma$-stable objects of class $v$ remain unchanged for any stability condition $\sigma$ in the same connected chamber;
@@
In general, the classification of walls is not available but one always has the following two situations for a stability condition.
@@
The walls and wall-crossing behaviors of the moduli space $\textsf{{M}}_{\sigma}(v)$ are well-understood for some stability conditions, on $\mathbb{P}^{2}$ [36], on K3 surfaces [6, 7, 39] and Enriques surfaces [40]. Since a strongly smooth Gushel–Mukai surface is in particular a K3 surface, the article uses some results regarding wall-crossing for K3 surfaces which are recollected in what follows.
@@ 3.3. Some stability conditions on K3 surfaces
$\mathcal{A}_{\omega,\beta}=\{E\in{\textbf{{D}}}^{b}(S)\,|\,\mathscr{H}^{i}(E)=0\textup{ for }i\notin\{-1,0\},\mathscr{H}^{-1}(E)\in\mathcal{F}_{\omega,\beta},\mathscr{H}^{0}(E)\in\mathcal{T}_{\omega,\beta}\}$
@@
Proposition 3.12 ([14, Lemma 6.2 and Section 7]).
@@
using the above criterion and Proposition 2.4. Otherwise, one has $Z_{\omega,\beta}(w)\leq 0$ for some spherical class $w=(r,\Delta,s)$ with $r>0$. In particular, one has
$r>0,\quad\Delta^{2}=2rs-2,\quad\Delta\cdot H=-4r,\quad 3r+2\Delta\cdot H+5s\geq 0$
at the same time. Set x=rx=-r and $y=\Delta^{2}/2$, then the only possible integer solutions for $(x,y)$ are (1,0)(-1,0) and (2,3)(-2,3). It follows that one has either $H\cdot\Delta=4$ and $\Delta^{2}=0$ or $H\cdot\Delta=2$ and $\Delta^{2}=0$. The first case is impossible due to Proposition 2.4, and the second one by very ampleness of $H$ and Reider’s method [49].
@@ 3.4. The walls and wall-crossing on K3 surfaces
Let $S$ be a K3 surface and $v\in\textsf{{H}}(S,\mathbb{Z})$ be a primitive Mukai vector with $\langle v,v\rangle\geq-2$. Suppose that $W\subset\operatorname{Stab}^{{\dagger}}(S)$ is a wall with respect to $v$ in the sense of Proposition 3.8 and $\sigma_{0}$ is a generic point on $W$, then for any two $v$-generic stability conditions $\sigma_{\pm}$ in the adjacent chambers with respect to $W$ one has a birational map
@@
and two contractions $\textsf{{M}}_{\sigma_{\pm}}(v)\rightarrow\bar{M}_{\pm}$ contracting $S$-equivalent objects by [6, Theorem 1.4].
@@
Proposition 3.15 ([7, Proposition 5.1]).
Under the setups in this subsection, the Mukai vector of any HN factor of an object $A$ in $\textsf{{M}}_{\sigma_{+}}(v)$ with respect to $\sigma_{-}$ is contained in $\textbf{{H}}_{W}$, and vice versa. The same holds for any object in $\textsf{{M}}_{\sigma_{0}}(v)$. Similarly, any Jordan–Hölder factor of an object $A\in\textsf{{M}}^{\textsf{{st}}}_{\sigma_{0}}(w)$ for some $w\in\textbf{{H}}_{W}$ will have Mukai vector contained in $\textbf{{H}}_{W}$.
In this case, using the existence of certain semistable objects, one can determine the type of the wall $W$ according to specific vectors in $\textbf{{H}}_{W}$ by [7, Theorem 5.7] when $\langle v,v\rangle>0$.
@@
Let $\langle v,v\rangle>0$, then the wall $W$ is called a fake wall once there are no curves in $\textsf{{M}}_{\sigma_{\pm}}(v)$ that are $S$-equivalent to each other with respect to $\sigma_{0}$; a flopping wall once $\bar{M}_{+}=\bar{M}_{-}$ and the induced birational map $\textsf{{M}}_{\sigma_{-}}(v)\dashrightarrow\textsf{{M}}_{\sigma_{+}}(v)$ induces a flopping contraction; a divisorial wall once the induced contractions $\textsf{{M}}_{\sigma_{\pm}}(v)\rightarrow\bar{M}_{\pm}$ are divisorial contractions.
@@
In this case, according to [7, Proposition 9.4 and Section 14], the Jordan–Hölder factors for a strictly semistable object $A$ in $\textsf{{M}}_{\sigma_{0}}(v)$ are spherical $\sigma_{\pm}$-stable objects. Moreover, one notices that the vector $w^{\prime}=v-w$ also satisfies $\langle w^{\prime},w^{\prime}\rangle=-2$ and $\langle v,w^{\prime}\rangle=1$ in the setting of Theorem 3.18, so a general flopping wall is induced by two spherical stable objects.
@@ 4. The passage to special Gushel–Mukai threefolds
4.1. Equivariant categories
Given an action of $G$ on $\mathcal{D}$, one can construct the equivariant category $\mathcal{D}_{G}$ with respect to the $G$-action on $\mathcal{D}$ whose objects and morphisms are respectively the $G$-equivariant objects and $G$-equivariant morphisms according to [22, 10].
Suppose in addition that $G$ is abelian, then there exists a natural action of the dual group $G^{\vee}:=\operatorname{Hom}(G,\mathbb{C}^{\times})$ on $\mathcal{D}_{G}$ induced by a $G$-action on $\mathcal{D}$.
Theorem 4.1 ([22, Theorem 4.2]).
Consider a finite abelian group $G$ acting on an idempotent complete category $\mathcal{D}$, then there exists an equivalence ${\textup{{{X}}}}\colon(\mathcal{D}_{G})_{G^{\vee}}\cong\mathcal{D}$ such that the $G$-action is identified with the $(G^{\vee})^{\vee}$-action by $(G^{\vee})^{\vee}\cong G$.
@@ 4.1. Equivariant categories
given on objects by sending $A$ to $\bigoplus_{g\in G}{\textup{{{A}}}}_{g}A$ with the natural $G$-equivariant structure (one can see, for example, [22, Lemma 3.8]). Moreover, one has the following observation:
@@
are exact. Then, according to Proposition 4.2, the equivalence X in Theorem 4.1 respects exact triangles and is therefore an equivalence of triangulated categories.
@@ 4.2. Equivariant stability conditions and covering maps
Proposition $4.$
In the setup of Theorem 4.3, the stability condition $(\sigma_{\mathfrak{S}_{2}})_{\mathfrak{S}_{2}^{\vee}}$ corresponds to the stability condition $2\cdot\sigma=(\mathcal{P},2Z)$ under the equivalence ${\textup{{{X}}}}\colon(\mathcal{D}_{\mathfrak{S}_{2}})_{\mathfrak{S}_{2}^{\vee}}\cong\mathcal{D}$ of Theorem 4.1.
@@ 4.3. The Kuznetsov components and the involutions
The numerical Grothendieck group $\textsf{{N}}({\textbf{{Ku}}}(X))$ has rank 2. Denote the first Chern class of the polarization $\mathscr{O}_{X}(1)$ by $h$, then one has a basis $\kappa_{1},\kappa_{2}$ of $\textsf{{N}}({\textbf{{Ku}}}(X))$ with
$\textup{{ch}}(\kappa_{1})=1-\frac{1}{5}h^{2}\quad\textup{and}\quad\textup{{ch}}(\kappa_{2})=2-h+\frac{1}{12}h^{3}$
@@
The opposite of $X$ is a strongly smooth Gushel–Mukai surface $S$ and one has a covering involution $f\colon X\rightarrow X$ for the double cover $X\rightarrow M_{S}$ branched over $S\subset M_{S}$. The direct image of the covering involution restricts to an involution on the Kuznetsov component ${\textbf{{Ku}}}(X)$ and induces a $\mathfrak{S}_{2}$-action on ${\textbf{{Ku}}}(X)$. According to [33, 47], the equivariant category ${\textbf{{Ku}}}(X)_{\mathfrak{S}_{2}}$ is equivalent to ${\textbf{{D}}}^{b}(S)$ and the dual $\mathfrak{S}_{2}$-group action on the equivariant category is induced by an involution on ${\textbf{{D}}}^{b}(S)$ say $\Pi$. The involution $\Pi$ induces a Hodge isometry on $\textsf{{H}}(S,\mathbb{Z})$ whose invariant lattice is spanned by $v_{1}=(1,0,-1)$ and $v_{2}=(-2,H,-2)$.
@@
$\textup{{inf}}(a\kappa_{1}+b\kappa_{2})$ $=-a(1,0,-1)+b(-2,H,-2)$
@@
Moreover, the stability condition $\sigma_{S}$ in Example 3.13 is invariant under the involution $\Pi$ according to [37, Proposition 6.2].
5. Double dual EPW sextics as Bridgeland moduli spaces